# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v4

@@ Appendix A LLM Prompts and Responses
LLM Generated Responses LLM Response of pz-mpe-simple-spread: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batchsize = agents_q.size(0) n_agents = 6 global_state_reshaped = global_state.view(batchsize, n_agents, 36) # Process landmark distances landmarks_relative = global_state_reshaped[:, :, 4:16].view(batchsize, n_agents, 6, 2) landmark_distances = torch.norm(landmarks_relative, dim=-1) min_dist_per_agent = torch.clamp(torch.min(landmark_distances, dim=-1)[0], min=0.1) pos_weights = 1.0 / min_dist_per_agent # Process collision counts positions = global_state_reshaped[:, :, 2:4] diffs = positions.unsqueeze(2) - positions.unsqueeze(1) pairwise_dist = torch.norm(diffs, dim=-1) collision_mask = (pairwise_dist <<= 0.3)  &  torch.eye(n_agents, device=positions.device).bool().unsqueeze(0) collision_weights = collision_mask.sum(dim=-1).float() # Combine weights and softmax combined_weights = pos_weights - collision_weights softmax_weights = torch.softmax(combined_weights, dim=-1) global_q = (agents_q  softmax_weights).sum(dim=-1, keepdim=True) return global_q LLM Response of lbforaging:Foraging-8x8-2p-2f-coop-v3: import torch def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) agents_data = global_state[:, :6].view(batch_size, 2, 3) food_data = global_state[:, 6:12].view(batch_size, 2, 3) agents_level = agents_data[:, :, 2] agents_pos = agents_data[:, :, :2].unsqueeze(2) food_pos = food_data[:, :, :2].unsqueeze(1) dxdy = torch.abs(agents_pos - food_pos).sum(dim=-1) adjacent = (dxdy == 1).float() active_food_mask = ((food_data[:, :, 0] != -1) & (food_data[:, :, 1] != -1)).unsqueeze(1) adjacent = adjacent  active_food_mask total_contribution = torch.zeros_like(agents_level) for j in range(2): adjacent_j = adjacent[:, :, j] food_level_j = food_data[:, j, 2] sum_adj_levels = (agents_level  adjacent_j).sum(dim=1) num_adj = adjacent_j.sum(dim=1) valid = ((num_adj >>= 2) & (sum_adj_levels >>= food_level_j)).float().unsqueeze(1) contribution_j = adjacent_j  food_level_j.unsqueeze(1)  valid total_contribution += contribution_j agent_contrib = agents_level  total_contribution weights = torch.softmax(agent_contrib, dim=1) global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of academy_counterattack_easy: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: batch_size = agents_q.size(0) n_agents = agents_q.size(1) # Extract ball information from first agent’s observation ball_pos = global_state[:, 88:90] holder_info = global_state[:, 94:97] is_my_team = holder_info[:, 1].bool() # Calculate distance to goal goal_pos = torch.tensor([1.0, 0.0], device=agents_q.device) ball_goal_dist = torch.norm(ball_pos - goal_pos, dim=1) valid_advance = is_my_team & (ball_goal_dist >> 0.19) & (ball_goal_dist << 0.99) # Calculate agent-ball distances and holder mask all_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-ball_pos[:,0], pos_y-ball_pos[:,1]], dim=1), dim=1) all_dists.append(dist.unsqueeze(1)) dist_matrix = torch.cat(all_dists, dim=1) holder_mask = F.one_hot(torch.argmin(dist_matrix, dim=1), num_classes=n_agents).bool() # Calculate agent-goal distances goal_dists = [] for i in range(n_agents): agent_obs = global_state[:, i115:(i+1)115] agent_id = torch.argmax(agent_obs[:, 97:108], dim=1) pos_x = agent_obs[torch.arange(batch_size), 2agent_id] pos_y = agent_obs[torch.arange(batch_size), 2agent_id+1] dist = torch.norm(torch.stack([pos_x-goal_pos[0], pos_y-goal_pos[1]], dim=1), dim=1) goal_dists.append(dist.unsqueeze(1)) goal_dist_matrix = torch.cat(goal_dists, dim=1) # Calculate weights components advance_component = holder_mask.float()  valid_advance.unsqueeze(1)  5.0 position_component = 1.0 / (goal_dist_matrix + 1e-6) # Combine components with softmax combined = torch.where(valid_advance.unsqueeze(1), advance_component, position_component) weights = F.softmax(combined, dim=1) # Compute global Q-value global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q LLM Response of matrixgames:climbing-nostate-v0: def QLLMNetwork(agents_q: torch.Tensor, global_state: torch.Tensor) ->> torch.Tensor: device = agents_q.device batch_size = agents_q.shape[0] # Initialize base weights and reward matrix weights = torch.full((batch_size, 2), 0.5, device=device) reward_matrix = torch.tensor([[11, -30, 0], [-30, 7, 0], [0, 6, 5]], device=device, dtype=torch.float32) # Create mask for valid states (non-initial) valid_mask = (global_state[:, 0] >>= 0) & (global_state[:, 1] >>= 0) if valid_mask.any(): prev_actions = global_state[valid_mask].long() a0 = prev_actions[:, 0] a1 = prev_actions[:, 1] # Calculate previous rewards and optimal condition prev_rewards = reward_matrix[a0, a1] optimal_mask = (prev_rewards == 11) # Calculate deviation from optimal for each agent agent0_deviation = (a0 != 0).float() agent1_deviation = (a1 != 0).float() # Create penalty weights based on deviation penalty_weights = torch.stack([1.0 - agent0_deviation  0.5,1.0 - agent1_deviation  0.5], dim=1) # Apply optimal condition override penalty_weights[optimal_mask] = 0.5 # Maintain equal weights for optimal case # Normalize weights to maintain sum=1 weight_sums = penalty_weights.sum(dim=1, keepdim=True) normalized_weights = penalty_weights / weight_sums # Update weights for valid samples weights[valid_mask] = normalized_weights # Compute global Q-value with dynamic weighting global_q = (agents_q  weights).sum(dim=1, keepdim=True) return global_q
Appendix B Details of Environments and Baselines
Baseline Algorithms
@@ Appendix B Details of Environments and Baselines
Code Repository
The baseline algorithm in this experiment is implemented based on the epymarl [^5] https://github.com/uoe-agents/epymarl and pymarl2666https://github.com/hijkzzz/pymarl2 codebases. We tested MPE, LBF, and matrix games using the epymarl library, while Google Research Football was evaluated with pymarl2. Although the SMAC [^7] https://github.com/oxwhirl/smac environment is a popular benchmark for multi-agent reinforcement learning and is built upon PySC2888https://github.com/google-deepmind/pysc2, it lacks comprehensive documentation regarding the specific unit types, abilities, and map configurations used in each scenario. This obscurity hinders reproducibility and limits fine-grained control over experimental variables. Consequently, we opted for several environments with better transparency and configurability to ensure experimental rigor and facilitate controlled evaluations.
Tasks Details
Matrix Games
The cooperative matrix games environment is a simple benchmark setting in multi-agent reinforcement learning, where two or more agents simultaneously select actions from predefined action spaces. The environment is defined by a payoff matrix that specifies the global reward received based on the joint actions of all agents (below is the payoff matrix used in test environment). Cooperation is essential, as the optimal strategy often depends on coordinated decision-making among agents. In order to make sense of the global state space of this task, we include the actions executed by the agents last step.
$\begin{bmatrix}0&6&5\\ -30&7&0\\ 11&-30&0\end{bmatrix}$
Figure 7: Environment visualizations of Level-Based Foraging (LBF), Google Research Football (GRF), and Multi-Agent Particle Environments (MPE)
Level-Based Foraging
@@
Multi-Agent Particle Environments
@@
Google Research Football
Table 3: Details of GRF tasks
@@
GRF tasks n_agents state shape action num
academy_3_vs_1_with_keeper 4 460 19
academy_counterattack_easy 4 460 19
academy_pass_and_shoot_with_keeper 3 345 19
Appendix C Additional Experiments
Compatibility Testing of Different LLMs
We tested the performance of QLLM with three mainstream LLMs, Deepseek-R1, Deepseek-V3999https://www.deepseek.com/, and ChatGPT-4o [^10] https://openai.com/, along with other algorithms in the MPE environment. As shown in Figure 8, the experimental results indicate that mainstream LLMs outperform the baseline algorithms. Moreover, as the inference capabilities of the LLMs improve, the performance of our algorithm also improves.
Figure 8: Performance comparison of QLLM using different LLMs in the GRF environment.
Tests of the Effect of the Evaluator LLM
Most reinforcement learning algorithms enhanced by LLMs use only one kind of LLM to do their job directly. Our coder-evaluator framework features two LLMs, each with its specific role, working together to generate an accurate credit assignment function. Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in LLMs. We tested the performance of QLLM across different reinforcement learning environments without using the evaluator LLM. The experimental results (Figure 9) show that, the evaluator LLM significantly improves the algorithm’s performance in most environments.
Figure 9: Performance comparison of QLLM without evaluator LLM in the GRF environment, where QLLM-C stands for QLLM that do not use the evaluator LLM
Comparison of Algorithms Training Costs
As mentioned in the main text, QLLM replaces the mixing network with a deterministic functional expression, TFCAF, resulting in a significantly smaller number of learnable parameters. Therefore, the training cost of our algorithm is reduced. Table 4 lists the number of trainable parameters for each algorithm in the environments tested in this paper. Our algorithm reduces the number of trainable parameters by approximately 12% to 37%. This reduction directly lowers the training cost and simplifies the model structure, making the algorithm more efficient and improving performance.
Table 4: Number of learnable parameters (in K) for each environment under different algorithms
Environment / Algorithm QLLM QMIX QATTEN MADDPG RIIT QPLEX Baseline Average Reduction Rate (%)
pz-mpe-simple-spread 105.221 140.422 191.114 154.246 191.175 154.762 166.744 36.88%
pz-mpe-simple-tag 102.277 112.806 129.034 127.494 132.199 128.010 125.909 18.78%
pz-mpe-simple-adversary 101.381 107.270 118.282 122.246 124.359 122.762 118.584 14.52%
Foraging-2s-15x15-4p-3f 103.174 111.815 121.963 126.727 130.536 127.372 123.283 16.30%
Foraging-2s-8x8-2p-2f-coop 101.766 106.887 116.683 122.119 124.296 122.764 118.150 13.87%
Foraging-2s-10x10-3p-3f 102.662 109.735 120.043 124.807 127.912 125.452 121.190 15.28%
climbing-nostate-v0 100.099 103.684 110.408 118.148 118.661 118.406 113.861 12.07%
academy_3_vs_1_with_keeper 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
academy_counterattack_easy 430.355 529.300 601.496 634.900 568.085 543.014 575.359 25.23%
academy_pass_and_shoot_with_keeper 430.099 504.884 560.760 600.084 530.869 523.046 543.129 20.83%

### v5

@@ Appendix 0.A LLM Prompts and Responses
Appendix 0.B Additional Experiments
0.B.1 Ablation Study on the Evaluator LLM
While most LLM-enhanced reinforcement learning methods rely on a single model, our coder-evaluator framework leverages the collaboration between $M_{\text{coder}}$ and $M_{\text{evaluator}}$ to ensure logical precision. To isolate the impact of the evaluator, we evaluated a variant termed QLLM-C, which utilizes only the coder LLM for single-shot generation (effectively $K=1,T=1$).
We conducted this ablation study on the SMAC 3s_vs_5z and 2c_vs_64zg scenarios. As illustrated in Figure 8, QLLM-C exhibits a slight performance advantage over the QMIX baseline, suggesting that the zero-shot prior knowledge encoded in LLMs can capture basic coordination patterns. However, QLLM-C is significantly outperformed by the full QLLM framework in both scenarios. This performance gap demonstrates that $M_{\text{evaluator}}$ is essential for filtering out hallucinations and refining credit assignment logic through iterative feedback, especially in complex environments where high-fidelity reasoning is critical. These results validate the architectural necessity of our coder-evaluator framework.
Figure 8: Performance comparison of QLLM and QLLM-C on SMAC maps. QLLM-C stands for the variant utilizing only the coder LLM for generation.
0.B.2 Ablation Study on Candidate Number and Iteration Rounds
To investigate the impact of the number of candidate TFCAFs per round ($K$) and the number of iteration rounds within the coder-evaluator framework ($T$), we conducted parameter ablation studies on the MPE simple-tag task. As illustrated in Figure 9, the results reveal that in relatively simple scenarios like MPE, generating a single TFCAF without iterative optimization (i.e., $K=1,T=1$) is already sufficient to achieve near-optimal performance.
However, this finding contrasts with our observations in complex environments such as SMAC. In fact, the QLLM-C variant (refer to Fig. 8) effectively represents the $K=1,T=1$ configuration, and its performance in SMAC scenarios is significantly inferior to that of the standard $QLLM(K=3,T=3)$. We can conclude that for low-complexity tasks, minimal generation counts and iterations are adequate. In contrast, complex scenarios necessitate a larger candidate pool and multiple refinement cycles to synthesize TFCAFs with sophisticated and robust coordination logic.
Figure 9: Ablation study of the number of candidates $K$ and iteration rounds $T$ on the pz-mpe-simple-tag scenario.
0.B.3 Compatibility Across Different LLMs
To assess framework compatibility, we implemented QLLM using DeepSeek-R1 [7], DeepSeek-V3111https://www.deepseek.com/, and Mistral-Large [^2] https://mistral.ai/ on the MPE simple-tag task. As illustrated in Fig. 10, QLLM consistently surpasses the QMIX baseline regardless of the underlying LLM architecture. Furthermore, the results show a positive correlation between LLM reasoning capabilities and RL performance; specifically, the Mistral-Large variant exhibits a slight performance margin over the others. These findings demonstrate that QLLM is a versatile framework capable of effectively leveraging the diverse reasoning strengths of various state-of-the-art LLMs.
Figure 10: Performance comparison of QLLM variants using different mainstream LLMs on the MPE simple-tag scenario.
0.B.4 Comparison of Algorithms Training Costs
As discussed in the methodology section, QLLM replaces the traditional neural mixing network with a deterministic functional expression, TFCAF. This architectural shift results in a significantly lower number of learnable parameters compared to standard value decomposition methods. Table 1 details the specific number of trainable parameters for each algorithm across the tested environments, while Figure 11 provides a visual comparison of the average parameter counts. As illustrated, QLLM reduces the number of trainable parameters by approximately 13% to 37% relative to the baseline average. This substantial reduction simplifies the model structure and directly lowers the training cost, thereby enhancing the overall efficiency of the algorithm while maintaining superior performance.
Furthermore, we evaluated the operational efficiency of QLLM by comparing its average training duration against QMIX across these SMAC scenarios. Experimental results show that QLLM reduces the average training time per step from 161.00 ms to 104.18 ms (35.3% reduction) and shortens the average total duration for 2 million training steps from approximately 9.05 hours to 5.38 hours (40.5% reduction). Although QLLM includes an additional phase for TFCAF synthesis, one iteration of the code generation process (with $n=3$) takes only about 15 minutes, which is negligible compared to the several hours of training time saved. These findings demonstrate that QLLM not only achieves superior performance but also significantly optimizes training costs and time efficiency by reducing learnable parameters and accelerating the training procedure.
Table 1: Number of learnable 0(in K) for each environment under different algorithms
Environment / Algorithm QLLM QMIX QPLEX Qatten RIIT Baseline Average Reduction Rate
pz-mpe-simple-spread 105.22 140.42 154.76 191.11 191.18 166.74 36.88%
pz-mpe-simple-tag 102.28 112.81 128.01 129.03 132.20 125.91 18.78%
pz-mpe-simple-adversary 101.38 107.27 122.76 118.28 124.36 118.58 14.52%
Foraging-2s-15x15-4p-3f 103.17 111.82 127.37 121.96 130.54 123.28 16.30%
Foraging-2s-8x8-2p-2f-coop 101.77 106.89 122.76 116.68 124.30 118.15 13.87%
Foraging-2s-10x10-3p-3f 102.66 109.74 125.45 120.04 127.91 121.19 15.28%
academy_3_vs_1_with_keeper 430.36 529.30 543.01 601.50 568.09 575.36 25.23%
academy_counterattack_easy 430.36 529.30 543.01 601.50 568.09 575.36 25.23%
academy_pass_and_shoot_with_keeper 430.10 504.88 523.05 560.76 530.87 543.13 20.83%
Figure 11: Comparison of the average number of learnable parameters across all tested environments.
0.B.5 Experimental Setup
The experiments were conducted on computing nodes equipped with an Intel(R) Xeon(R) Platinum 8488C CPU and NVIDIA H100 PCIe GPUs with 80GB of VRAM. Our framework was implemented using Python 3.8.20, with PyTorch 2.1.0 serving as the deep learning backend. To ensure high-performance GPU acceleration, the system was configured with NVIDIA Driver version 570.195.03 and CUDA 12.8.
Appendix 0.C Details of Environments and Baselines
0.C.1 Baseline Algorithms
@@
0.C.2 Code Repository
The baseline algorithms in this experiment are implemented based on the epymarl [^3] https://github.com/uoe-agents/epymarl, pymarl [^4] https://github.com/oxwhirl/pymarl, and pymarl2555https://github.com/hijkzzz/pymarl2 codebases. Specifically, we evaluated the MPE and LBF environments using the epymarl library. The StarCraft Multi-Agent Challenge (SMAC) was tested utilizing the original pymarl framework, while Google Research Football (GRF) was evaluated with pymarl2. This setup ensures experimental rigor and facilitates fair and controlled comparisons across diverse cooperative tasks.
0.C.3 Tasks Details
Figure 12: Environment visualizations of Level-Based Foraging (LBF), Google Research Football (GRF), Multi-Agent Particle Environments (MPE) and StarCraft Multi-Agent Challenge (SMAC).
0.C.3.1 Level-Based Foraging
@@
0.C.3.2 Google Research Football
Table 3: Details of GRF tasks
GRF Tasks N_agents State Shape Action Num
academy_3_vs_1_with_keeper 4 460 19
academy_counterattack_easy 4 460 19
academy_pass_and_shoot_with_keeper 3 345 19
0.C.3.3 Multi-Agent Particle Environments
The Multi-Agent Particle Environments are a widely-used set of communication-oriented environments where particle agents can move, communicate, observe each other, exert physical forces, and interact with fixed landmarks within a 2D continuous space[1]. These environments are highly configurable and serve as standard benchmarks for evaluating coordination, competition, and communication in MARL.
@@ 0.C.3.3 Multi-Agent Particle Environments
0.C.3.4 StarCraft Multi-Agent Challenge
The StarCraft Multi-Agent Challenge is a high-dimensional micro-management benchmark based on the StarCraft II engine. In this environment, agents must learn to cooperate with each other and execute sophisticated combat strategies (such as focus fire and kiting) to defeat enemy units and achieve victory. To evaluate the performance of our algorithm in complex multi-agent cooperative scenarios, we selected two challenging maps categorized as hard difficulty: 3s_vs_5z and 2c_vs_64zg. These scenarios require sophisticated micromanagement and coordination to handle heterogeneous units or large-scale enemy swarms. The specific parameters of these two maps are detailed in Table 5.
Table 5: Details of the SMAC scenarios.
@@ 0.C.3.4 StarCraft Multi-Agent Challenge
SMAC Tasks Ally Units Enemy Units State Shape Action Num
3s_vs_5z 3 Stalkers 5 Zealots 68 11
2c_vs_64zg 2 Colossi 64 Zerglings 342 70
Appendix 0.D Proof of the QLLM Decomposition Formula
In this section, we provide a formal derivation of the decomposition formulation used in QLLM (Equation (3)). We aim to prove that the joint action-value function $Q_{\text{tot}}(s,\boldsymbol{a})$ can be represented as a state-dependent linear combination of individual utility functions $Q_{i}(\tau^{i},a^{i})$ plus a state-dependent bias term $f_{b}(s)$.
Assuming the joint value function $Q_{\text{tot}}$ is differentiable with respect to the individual $Q$-values, we can perform a multivariate Taylor expansion of $Q_{\text{tot}}$ at the state $s$ as a function of the utility vector $\boldsymbol{Q}=(Q_{1},\dots,Q_{n})$:
$Q_{\text{tot}}(s,\boldsymbol{Q})=\text{const}(s)+\sum_{i=1}^{n}\mu_{i}(s)Q_{i}+\sum_{i,j}\mu_{ij}(s)Q_{i}Q_{j}+\mathcal{O}(Q^{3}),$ (10)
where
$\mu_{i}(s)=\frac{\partial Q_{\text{tot}}}{\partial Q_{i}}\quad\text{and}\quad\mu_{ij}(s)=\frac{1}{2}\frac{\partial^{2}Q_{\text{tot}}}{\partial Q_{i}\partial Q_{j}}.$
To analyze the local behavior near an optimal joint action $\boldsymbol{a}_{0}=(a_{0}^{1},\dots,a_{0}^{n})$, we note that the gradient $\nabla_{\boldsymbol{a}}Q_{\text{tot}}$ vanishes at $\boldsymbol{a}_{0}$. Since
$\frac{\partial Q_{\text{tot}}}{\partial a^{i}}=\frac{\partial Q_{\text{tot}}}{\partial Q_{i}}\frac{\partial Q_{i}}{\partial a^{i}}=0,$
and assuming that each agent has a non-trivial impact on the team ($\mu_{i}\neq 0$), the local gradient of each agent’s utility must also vanish:
$\frac{\partial Q_{i}}{\partial a^{i}}(a_{0}^{i})=0.$
Consequently, each individual utility function can be approximated by its local quadratic expansion:
$Q_{i}(a^{i})\approx\alpha_{i}(s)+\beta_{i}(s)(a^{i}-a_{0}^{i})^{2}.$
The key to linearizing the formulation lies in the treatment of the second-order cross terms $Q_{i}Q_{j}$. Substituting the quadratic approximation into these terms yields:
$Q_{i}Q_{j}$ $\approx[\alpha_{i}+\beta_{i}(a^{i}-a_{0}^{i})^{2}][\alpha_{j}+\beta_{j}(a^{j}-a_{0}^{j})^{2}]$
$=\alpha_{i}\alpha_{j}+\alpha_{i}\beta_{j}(a^{j}-a_{0}^{j})^{2}+\alpha_{j}\beta_{i}(a^{i}-a_{0}^{i})^{2}+\textit{higher-order terms}$
$\approx\alpha_{i}\alpha_{j}+\alpha_{i}(Q_{j}-\alpha_{j})+\alpha_{j}(Q_{i}-\alpha_{i})$
$=\alpha_{j}Q_{i}+\alpha_{i}Q_{j}-\alpha_{i}\alpha_{j}.$ (11)
Equation (11) demonstrates that the product of two utilities near the optimum can be effectively represented as a linear combination of $Q_{i}$ and $Q_{j}$ plus a state-dependent constant. By substituting this result into the Taylor expansion in Equation (10) and grouping all terms associated with each $Q_{i}$, we obtain:
$\begin{split}Q_{\text{tot}}(s,\boldsymbol{a})&\approx\sum_{i=1}^{n}\underbrace{\left[\mu_{i}(s)+2\sum_{j\neq i}\mu_{ij}(s)\alpha_{j}(s)+\dots\right]}_{f_{w}^{i}(s)}Q_{i}\\ &\quad+\underbrace{\left[\text{const}(s)-\sum_{i,j}\mu_{ij}(s)\alpha_{i}\alpha_{j}+\dots\right]}_{f_{b}(s)}.\end{split}$ (12)
This derivation confirms that the formulation $Q_{\text{tot}}(s,\boldsymbol{a})=\sum f_{w}^{i}(s)Q_{i}+f_{b}(s)$ is a mathematically sound approximation for any complex value-based mixing mechanism that satisfies the Individual-Global-Max (IGM) principle. While traditional methods like QMIX or Qatten utilize hypernetworks to estimate these coefficients, QLLM leverages the logical reasoning of LLMs to generate the functions $f_{w}^{i}(s)$ and $f_{b}(s)$ directly, which enhances both interpretability and parameter efficiency.